OpenStudy (anonymous):

f(x)=sqrt(x-3) g(x)=sqrt(X^2+3) find f o g and g o f

OpenStudy (jamesj):

(f o g)(x) = f(g(x)). So evaluate the expressions above that way. f(g(x)) = f(sqrt(x^2+3)) = ..... what?

OpenStudy (jamesj):

For example, if h(x) = x^2 and j(x) = x+1, then (h o j)(x) = h(j(x)) = h(x+1) = (x+1)^2

OpenStudy (anonymous):

i just wanted to check my answers i got fog = \[\sqrt{\sqrt{x^2+3}-3}\] and gof=\[\sqrt{x}\]

OpenStudy (matt101):

\[f(g(x)) = \sqrt{\sqrt{x^2+3}-3}\] \[g(f(x))=\sqrt{(\sqrt{x-3})^2+3}\]\[= \sqrt{x}\]

OpenStudy (matt101):

Yes that's correct